Ford07:

COULD I PLEASE HAVE HELP? Rewrite the expression with rational exponents as a radical expression by extending the properties of integer exponents. (1 point) y to the one third power all over y to the one sixth power the square root of y to the one sixth power the ninth root of y squared the square root of y to the sixth power the sixth root of y

Vocaloid:

\[\frac{ y^{\frac{ 1 }{ 3 }} }{ y^{\frac{ 1 }{ 6}} }\] when you have two of the same base divided like this, you can keep the base while subtracting the exponents. in this case, it would simply be y ^ (1/3 - 1/6). from there, use the properties of fractional exponents to re-write the expression as a root. recall: \[x^{\frac{ b }{ a }} =\sqrt[a]{x^{b}}\] notice how the numerator b becomes the exponent inside the radicand and the denominator a becomes the root.